When methane is burned with oxygen, the products are carbon dioxide and water. If you produce 9 grams of water and 11 grams of carbon dioxide from 16 grams of oxygen, how many grams of methane were needed for the reaction?
1
Expert's answer
2013-01-28T10:16:27-0500
First you need to write a balanced chemical equation. You are given thatmethane is burned, meaning a combustion reaction in which carbon dioxide and water are released.

Unbalanced: CH4 + O2 ---> CO2 + H2O
Balanced: CH4 + 2O2 ---> CO2 + 2H2O

Givens:
X grams CH4 (Molecular mass 16.0 grams)
9 grams H2O (Molecular mass 18.0 grams)
11 grams CO2 (Molecular mass 44.0 grams)
Mole ratio: 1:2:1:2 (CH4:O2:CO2:H2O)

Then you need to find which of the reactants are the limiting reactant and which is the excess reactant. The limiting reactant is what you will base the rest of the problem on. To do this, you convert each measurement to moles from grams.

11 g CO2 / (44.0 g) = 0.25 moles CO2

9 g H2O / (18.0 g) = 0.5 moles

n of CH4 = n of CO2 = n of H2O /2 = 0.25 moles

m of CH4 = n* Mw = 0.25 * 16.0 = 4 g